# Accurate and Efficient Low-Rank Model Merging in Core Space

### v1

@@ 4 The Core Space Merging Framework
Figure 2: Full Space Merging (left) firstly reconstructs full space matrices $\Delta W^{(t)}=B^{(t)}A^{(t)}$, and then performs merging in the full space to obtain $\Delta W$. KnOTS Merging concatenates the $\Delta W^{(t)}$ matrices, and performs a costly SVD on this high-dimensional matrix. Then, the $V^{(t)}$ matrices are merged and used to obtain the final $\Delta W$. The proposed Core Space Merging (right) performs SVD on a concatenation of low-dimensional $A^{(t)}$ and $B^{(t)}$ matrices to obtain reference bases. Afterwards, it performs SVD on the individual $A^{(t)}$ and $B^{(t)}$ matrices and calculates the core matrices. It then performs merging in the Core Space and reconstructs to obtain the final $\Delta W$.
@@ 4.1 Model Merging in Core Space
Let $A^{(t)}\in\mathbb{R}^{r\times n}$ and $B^{(t)}\in\mathbb{R}^{m\times r}$ denote the low-rank matrices for task $t$, derived from a shared pre-trained base model $W_{0}$. Each task update $\Delta W^{(t)}=B^{(t)}A^{(t)}$ can be reconstructed from the reduced SVD [^1] We use the reduced SVD throughout, as it offers a suitable alternative for low-rank matrices where $r\ll m,n$. of the matrices:
$A^{(t)}=U_{A}^{(t)}\Sigma_{A}^{(t)}V_{A}^{(t)\top}\!\!\!\!,\,\,\,B^{(t)}=U_{B}^{(t)}\Sigma_{B}^{(t)}V_{B}^{(t)\top}\!\!\!\!,\,\,\,\Delta W^{(t)}=U_{B}^{(t)}\Sigma_{B}^{(t)}V_{B}^{(t)\top}U_{A}^{(t)}\Sigma_{A}^{(t)}V_{A}^{(t)\top}.$ (1)
The shapes of the matrices in the decomposition are: $U_{A}^{(t)}\in\mathbb{R}^{r\times r}$, $\Sigma_{A}^{(t)}\in\mathbb{R}^{r\times r}$, $V_{A}^{(t)}\in\mathbb{R}^{n\times r}$, $U_{B}^{(t)}\in\mathbb{R}^{m\times r}$, $\Sigma_{B}^{(t)}\in\mathbb{R}^{r\times r}$, and $V_{B}^{(t)}\in\mathbb{R}^{r\times r}$.
Definition 1 (Core Matrix).
Given the SVD of low-rank matrices $A^{(t)}\in\mathbb{R}^{r\times n}$ and $B^{(t)}\in\mathbb{R}^{m\times r}$, the core matrix $M^{(t)}\in\mathbb{R}^{r\times r}$ is defined as:
$M^{(t)}:=\Sigma_{B}^{(t)}V_{B}^{(t)\top}U_{A}^{(t)}\Sigma_{A}^{(t)}.$ (2)
Intuitively, the core matrix $M^{(t)}$ is a compact representation summarizing the directional transformation applied by the low-rank update of task $t$, expressed in the singular spaces of $A^{(t)}$ and $B^{(t)}$. We confirm this intuition in Sec.˜5.2.
Under the hypothesis that all tasks share approximately the same common bases $(U_{B},V_{A})$ such that $U_{B}\approx U_{B}^{(t)},V_{A}\approx V_{A}^{(t)},\,\forall t\in\{1,\dots,T\}$, the following holds:
$\Delta W=\sum_{t=1}^{T}\Delta W^{(t)}\approx U_{B}\left(\sum_{t=1}^{T}M^{(t)}\right)V_{A}^{\top},$ (3)
suggesting that, under aligned bases, the sum of low-rank updates (i.e., Task Arithmetic ilharco2023task ) can be approximated by merging the corresponding core matrices – reducing the merging problem to operations in a much smaller $r\times r$ space. However, task-specific bases are misaligned in practice, making it impossible to directly compare or merge task-specific core matrices as in Eq.˜3. Therefore, we aim to find a shared basis that can represent all tasks without loss of information. Intuitively, such a shared basis should span the subspace formed by the union of the individual task subspaces.
@@
Given a collection of low-rank matrices $\{A^{(t)},B^{(t)}\}_{t=1}^{T}$, we define as reference bases the orthonormal matrices $(U_{B}^{\mathrm{ref}},V_{A}^{\mathrm{ref}})$ obtained by performing SVD over the vertically stacked $A^{(t)}$ and horizontally stacked $B^{(t)}$ low-rank matrices across tasks.
@@
For each task $t$, we express the task-specific bases $(U_{B}^{(t)},V_{A}^{(t)})$ in the coordinate system defined by the reference bases $(U_{B}^{\mathrm{ref}},V_{A}^{\mathrm{ref}})$ by solving the following least-squares problems:
@@
$R_{B}^{(t)}={U_{B}^{\mathrm{ref}}}^{\top}U_{B}^{(t)},\quad Q^{(t)}_{A}={V_{A}^{\mathrm{ref}}}^{\top}V_{A}^{(t)}$ (6)
The aligned core matrix $\tilde{M}^{(t)}$ for task $t$ is then computed as:
$\tilde{M}^{(t)}=R_{B}^{(t)}M^{(t)}{Q_{A}^{(t)}}^{\top}.$ (7)
Finally, we merge the set of aligned core matrices $\{\tilde{M}^{(t)}\}_{t=1}^{T}$ using a function $\mathcal{M}$, yielding $M_{\text{merged}}=\mathcal{M}(\{\tilde{M}^{(t)}\}_{t=1}^{T})$, where $\mathcal{M}$ denotes any existing merging strategy. We then reconstruct the final weight matrix from $M_{\text{merged}}$ by projecting it back to the full space using the reference bases:
@@
Core Space merging offers several advantages over merging in the full space:
  • No loss of information – transformation to core space and back to full space is lossless (see Sec.˜4.2 for the proof of zero alignment error). Moreover, when the merging function is linear (e.g., Task Arithmetic), the model merged in full space will be the same as the model merged in core space $\tilde{M}^{(t)}$ and then reconstructed using the reference bases.
  • Efficiency – aligned core matrices $\tilde{M}^{(t)}\in\mathbb{R}^{T\cdot r\times T\cdot r}$ have considerably lower dimensionality than full space matrices $\Delta W^{(t)}\in\mathbb{R}^{m\times n}$ since $T,r\ll m,n$. Therefore, we can perform an expensive merging technique in the subspace of reduced dimensionality for a fraction of the total computational cost (see Sec.˜4.3 for detailed complexity analysis).
  • Efficacy – in Sec.˜5.1 we show that merging in Core Space yields better performance than merging in alternative spaces, when the merging function is non-linear. In Sec.˜5.2, we shed light on the reasons for Core Space superiority. Our approach is found to yield compact representations of each task and improve alignment between subspaces of each task.
@@ 4.2 No Information Loss in Core Space Representation
For each task $t$, reconstructing from the aligned core matrix $\tilde{M}^{(t)}$ exactly recovers the original low-rank update, i.e.:
$\Delta W^{(t)}=U_{B}^{(t)}M^{(t)}{V_{A}^{(t)}}^{\top}={U_{B}^{\mathrm{ref}}}\tilde{M}(t){V_{A}^{\mathrm{ref}}}^{\top}=\left(U_{B}^{\mathrm{ref}}{R_{B}^{(t)}}\right)M^{(t)}\left(V_{A}^{\mathrm{ref}}{Q^{(t)}_{A}}\right)^{\top}$. (9)
The outer expressions are given by construction (see Eqs.˜1, 2 and 7). The middle equality holds if $U_{B}^{(t)}=U_{B}^{\mathrm{ref}}{R_{B}^{(t)}}$ and $V_{A}^{(t)}=V_{A}^{\mathrm{ref}}{Q^{(t)}_{A}}$ or, equivalently, if the solutions of the least squares problems (defined in Eq.˜5) incur zero alignment error. Indeed, we show that the reference bases $U_{B}^{\mathrm{ref}}$ and $V_{A}^{\mathrm{ref}}$, computed via the SVD of the stacked matrices $B^{(t)}$ and $A^{(t)}$ (see Eq.˜4), minimize the total alignment error across all $T$ tasks, achieving an error of exactly zero. To illustrate this, we first analyze the alignment error for a single task $t$, explicitly focusing on $U_{B}^{\mathrm{ref}}$. Analogous conclusions hold symmetrically for $V_{A}^{\mathrm{ref}}$. For clarity, in the following derivations, we assume that $T\cdot r\leq m$ and $T\cdot r\leq n$, so that the total LoRA rank does not exceed the maximum possible rank of the target weight matrix. In Sec.˜A.4, we provide a more general formulation covering the case where this assumption is violated and show that the same zero alignment error result continues to hold.
@@
$\max_{U\in\mathcal{S}}\left\|{U_{B}^{(t)}}^{\top}U\right\|_{F}^{2}=\max_{U\in\mathcal{S}}\mathrm{tr}\left(U^{\top}{U_{B}^{(t)}}{U_{B}^{(t)}}^{\top}U\right),\quad\mathcal{S}=\left\{U\in\mathbb{R}^{m\times Tr}\;\middle|\;U^{\top}U=I_{T\cdot r}\right\},$ (11)
where $\mathrm{tr}(\cdot)$ denotes the trace operator. The optimization domain is restricted to the Stiefel manifold $\mathcal{S}$ (i.e., the set of matrices with orthonormal columns). The following lemma characterizes the solution to this optimization problem:
@@
A solution $U^{*}$ to the quadratic program in Eq.˜11 is given by a basis whose columns include the $r$ eigenvectors corresponding to nonzero eigenvalues of $B^{(t)}{B^{(t)}}^{\top}\in\mathbb{R}^{m\times m}$ or, equivalently, by the $r$ left singular vectors of the matrix $B^{(t)}$. Moreover, at the optimum, the objective attains its maximum value $r$, resulting in zero alignment error in Eq.˜10.
We refer the reader to Appendix A.3 for a detailed proof. Briefly, the result follows by applying the method of Lagrange multipliers to augment the optimization objective with the Stiefel manifold constraint and then enforcing stationarity by setting the gradient of the Lagrangian to zero.
Extension to multiple tasks. Achieving zero reconstruction error for a single model $t$ does not guarantee optimality for any other model $t^{\prime}\neq t$. Therefore, we aim to identify a reference basis $U^{*}$ that jointly optimizes Eq.˜11 across all $T$ models. We formulate this global problem as:
where $\mathbf{U}_{B}=\big[U_{B}^{(1)}\;\;U_{B}^{(2)}\;\;\dots\;\;U_{B}^{(T)}\big]$ denotes the horizontal concatenation of all $U_{B}^{(t)}$ matrices. The equality in Eq.˜12 follows directly from the linearity of the trace operator and the distributivity of matrix multiplication concerning matrix addition: $M^{\top}A_{1}M+M^{\top}A_{2}M=M^{\top}(A_{1}+A_{2})M$.
Exact merging via linear functions in Core Space. Thanks to the lossless reconstruction of each task-specific update (see Eq. (9)), if the merging function $\mathcal{M}$ is linear (such as in Task-Arithmetic), then merging in Core Space produces the same merged matrix as merging in full space. Specifically:
$\mathcal{M}(\{\Delta W^{(t)}\}_{t=1}^{T})=\mathcal{M}(\{{U_{B}^{\mathrm{ref}}}\tilde{M}(t){V_{A}^{\mathrm{ref}}}^{\top}\}_{t=1}^{T})=U_{B}^{\mathrm{ref}}\mathcal{M}(\{\tilde{M}(t)\}_{t=1}^{T}){V_{A}^{\mathrm{ref}}}^{\top}.$ (13)
It is worth noting that Core Space merging is computationally efficient and does not increase time complexity – even when task arithmetic is applied in Core Space – due to its lower dimensionality, as we will show in the next section. Moreover, when the merging function is non-linear, we will show in Sec.˜5.1 that it can improve performance compared to standard merging strategies applied in the full space at a fraction of the computational cost. For example, when dealing with Llama 3 8B, we have a speedup of more than $600\times$ with respect to KnOTS stoica2024knots for the Iso-C marczak2025task method.
@@
3:Compute SVDs: $\text{stack}(A^{(t)})=U_{A}\Sigma_{A}\underline{V_{A}^{\mathrm{ref}\top}}$, $\text{stack}(B^{(t)})=\underline{U_{B}^{\mathrm{ref}}}\Sigma_{B}V_{B}^{\top}$ $\triangleright$ reference bases
4:for $t=1$ to $T$ do
5:  Compute: $\tilde{M}^{(t)}=U_{B}^{\mathrm{ref}}B^{(t)}A^{(t)}V_{A}^{\mathrm{ref}\top}$ $\triangleright$ Eq.˜6 and Eq.˜7
6:Merge aligned core matrices: $M_{\mathrm{merged}}=\mathcal{M}(\{\tilde{M}^{(t)}\}_{t=1}^{T})$
7:return $\Delta W=U_{B}^{\mathrm{ref}}M_{\mathrm{merged}}V_{A}^{\mathrm{ref}\top}$ $\triangleright$ reconstructed merged model
Practical Implementation. While we introduce the per-task SVDs of $A^{(t)}$ and $B^{(t)}$ for theoretical clarity, in practice the core matrix can be computed more directly as $\tilde{M}^{(t)}\;=\;U_{B}^{\mathrm{ref}^{\top}}B^{(t)}A^{(t)}V_{A}^{\mathrm{ref}},$ without explicitly decomposing each LoRA matrix. This follows directly by substituting the definitions of $R_{B}^{(t)}$ and $Q_{A}^{(t)}$ from Eq.˜6 into Eq.˜7, and by recalling the definitions of $B^{(t)}$, $A^{(t)}$, and $M^{(t)}$ (see Eqs.˜1 and 2). This yields identical results and is marginally more computationally efficient. The decomposition of $A^{(t)}$ and $B^{(t)}$ helps formalize the projection process and support the derivations in Appendix˜A, including orthogonality, rank preservation, and the zero-error reconstruction proofs. For this reason, we retain the SVD-based formulation in the theoretical exposition. We report in Algorithm˜1 the algorithm for Core Space merging with this computationally optimized approach and the illustration in Sec.˜D.1.

### v3

@@ 4 The Core Space Merging Framework
Figure 2: Full Space Merging (left) firstly reconstructs full space matrices $\Delta W^{(t)}=B^{(t)}A^{(t)}$, and then performs merging in the full space to obtain $\Delta W$. KnOTS Merging concatenates the $\Delta W^{(t)}$ matrices, and performs a costly SVD on this high-dimensional matrix. Then, the $V^{(t)}$ matrices are merged and used to obtain the final $\Delta W$. The proposed Core Space Merging (right) performs SVD on a concatenation of low-dimensional $A^{(t)}$ and $B^{(t)}$ matrices to obtain reference bases $(V_{A}^{\mathrm{ref}},U_{B}^{\mathrm{ref}})$. Afterwards, it projects each update into the Core Space to obtain the core matrices $\{M^{(t)}\}_{t=1}^{T}$. It then performs merging in the Core Space and reconstructs to obtain the final $\Delta W$.
@@ 4.1 Model Merging in Core Space
Let $A^{(t)}\in\mathbb{R}^{r\times n}$ and $B^{(t)}\in\mathbb{R}^{m\times r}$ denote the low-rank matrices for task $t$, derived from a shared pre-trained base model $W_{0}$. Each task update $\Delta W^{(t)}=B^{(t)}A^{(t)}$ can be reconstructed from the SVD of the matrices:
$\begin{gathered}A^{(t)}=U_{A}^{(t)}\Sigma_{A}^{(t)}V_{A}^{(t)\top},\quad B^{(t)}=U_{B}^{(t)}\Sigma_{B}^{(t)}V_{B}^{(t)\top},\\[3.00003pt] \Delta W^{(t)}=U_{B}^{(t)}\Sigma_{B}^{(t)}V_{B}^{(t)\top}U_{A}^{(t)}\Sigma_{A}^{(t)}V_{A}^{(t)\top},\end{gathered}$ (1)
where the shapes of the matrices in the decomposition are: $U_{A}^{(t)}\in\mathbb{R}^{r\times r}$, $\Sigma_{A}^{(t)}\in\mathbb{R}^{r\times r}$, $V_{A}^{(t)}\in\mathbb{R}^{n\times r}$, $U_{B}^{(t)}\in\mathbb{R}^{m\times r}$, $\Sigma_{B}^{(t)}\in\mathbb{R}^{r\times r}$, and $V_{B}^{(t)}\in\mathbb{R}^{r\times r}$.
Motivation. Under the hypothesis that all tasks share approximately the same common bases $(U_{B},V_{A})$ such that $\forall t\in\{1,\dots,T\},\;U_{B}\approx U_{B}^{(t)}$ and $V_{A}\approx V_{A}^{(t)}$, we have:
$\Delta W=\sum_{t=1}^{T}\Delta W^{(t)}=\sum_{t=1}^{T}B^{(t)}A^{(t)}\approx U_{B}\left(\sum_{t=1}^{T}\Sigma_{B}^{(t)}V_{B}^{(t)\top}U_{A}^{(t)}\Sigma_{A}^{(t)}\right)V_{A}^{\top},$ (2)
where $\Sigma_{B}^{(t)}V_{B}^{(t)\top}U_{A}^{(t)}\Sigma_{A}^{(t)}\in\mathbb{R}^{r\times r}$ encodes the directional transformation applied by the low-rank update of task $t$. This suggests that, under aligned bases, the sum of low-rank updates (i.e., Task Arithmetic ilharco2023task ) can be reduced to merging operations in a much smaller $r\times r$ space.
Projecting into the Core Space. In practice, task-specific bases are not aligned, making direct merging of core matrices as in Eq.˜2 infeasible. Therefore, we aim to find a shared basis that can represent all tasks without loss of information and that enables merging to be performed in a reduced space compared to the full space $\Delta W^{(t)}$. Intuitively, such a shared basis should span the subspace formed by the union of the individual task subspaces.
@@
Given a collection of low-rank matrices $\{A^{(t)},B^{(t)}\}_{t=1}^{T}$, we define as reference bases the orthonormal matrices $U_{B}^{\mathrm{ref}}\in\mathbb{R}^{m\times Tr}$ and $V_{A}^{\mathrm{ref}}\in\mathbb{R}^{n\times Tr}$, obtained by performing SVD over the vertically stacked $A^{(t)}$ and horizontally stacked $B^{(t)}$ low-rank matrices across tasks:
@@
Although the reference bases $(U_{B}^{\mathrm{ref}},V_{A}^{\mathrm{ref}})$ span all task-specific directions, each task $t$ is originally expressed in its own local bases $(U_{B}^{(t)},V_{A}^{(t)})$. To express each update in the common coordinate system for merging, we solve the following least-squares problems:
@@
As we will show in Sec.˜4.2, $\|U_{B}^{\mathrm{ref}}R_{B}^{(t)}-U_{B}^{(t)}\|_{F}^{2}=0$, which allows us to substitute $U_{B}^{(t)}$ with $U_{B}^{\mathrm{ref}}R_{B}^{(t)}$, and similarly $V_{A}^{(t)}$ with $V_{A}^{\mathrm{ref}}Q_{A}^{(t)}$, in Eq.˜1, yielding:
$\Delta W^{(t)}=U_{B}^{\mathrm{ref}}\,\underbrace{R_{B}^{(t)}\Sigma_{B}^{(t)}V_{B}^{(t)\top}U_{A}^{(t)}\Sigma_{A}^{(t)}{Q_{A}^{(t)}}^{\top}}_{\text{task-$t$ update in reference coordinates}}{V_{A}^{\mathrm{ref}}}^{\top}.$ (6)
By substituting the least-squares solutions from Eq.˜5 and using the definitions of $B^{(t)}$ and $A^{(t)}$ from Eq.˜1, we can equivalently write:
$\Delta W^{(t)}=U_{B}^{\mathrm{ref}}\left({U_{B}^{\mathrm{ref}}}^{\top}B^{(t)}A^{(t)}{V_{A}^{\mathrm{ref}}}\right){V_{A}^{\mathrm{ref}}}^{\top}.$ (7)
This reformulation expresses each $\Delta W^{(t)}$ in the reference basis, enabling all updates to be compared or merged within a shared coordinate system.
Definition 2 (Core Matrix).
We define the core matrix $M^{(t)}$ as:
$M^{(t)}=\left({U_{B}^{\mathrm{ref}}}^{\top}B^{(t)}\right)\left(A^{(t)}V_{A}^{\mathrm{ref}}\right)\in\mathbb{R}^{Tr\times Tr}.$ (8)
This formulation generalizes the middle expression in Eq.˜2, where aligned task-specific bases were implicitly assumed. In contrast, the core matrix $M^{(t)}$ is expressed in the reference bases $(U_{B}^{\mathrm{ref}},V_{A}^{\mathrm{ref}})$ and thus does not rely on any alignment assumption. It encodes the directional transformation applied by the low-rank update of task $t$, providing a compact and lossless representation of each task update in the shared reference space and enabling efficient merging in a reduced $Tr\times Tr$ space.
Reparametrized Model Merging in Core Space. Once task-specific updates $\Delta W^{(t)}$ have been reparameterized into their corresponding core matrices $M^{(t)}$ in the shared reference bases $(U_{B}^{\mathrm{ref}},V_{A}^{\mathrm{ref}})$, Core Space Merging enables model merging to be performed entirely within a compact, aligned, low-rank subspace. Specifically, the merged update is computed by applying a merging operator $\mathcal{M}$ over the set of core matrices:
$M_{\text{merged}}=\mathcal{M}\big(\{M^{(t)}\}_{t=1}^{T}\big),$ (9)
where $\mathcal{M}$ may be any merging function, ranging from simple arithmetic averaging ilharco2023task to more advanced, non-linear or geometry-aware techniques yadav2023tiesmerging ; yu2024language . The final update in the original model space, as described also in Algorithm˜1, is recovered by projecting $M_{\text{merged}}$ back through the reference bases:
@@
Because Core Space is a lossless representation of the original updates for each individual task (see Eq.˜7), merging in this space preserves all relevant task information. Furthermore, when $\mathcal{M}$ is linear, such as Task Arithmetic, the merge operation in Core Space is exactly equivalent to applying the same merge in the full model space:
$\mathcal{M}(\{\Delta W^{(t)}\}_{t=1}^{T})=\mathcal{M}(\{{U_{B}^{\mathrm{ref}}}M^{(t)}{V_{A}^{\mathrm{ref}}}^{\top}\}_{t=1}^{T})=U_{B}^{\mathrm{ref}}\mathcal{M}(\{M^{(t)}\}_{t=1}^{T}){V_{A}^{\mathrm{ref}}}^{\top}.$ (11)
Core Space merging offers key benefits over full space merging:
  • Efficiency. Core matrices $M^{(t)}\in\mathbb{R}^{Tr\times Tr}$ are significantly smaller than their full-space counterparts $\Delta W^{(t)}\in\mathbb{R}^{m\times n}$. This reduction enables high-cost merging algorithms to run at a fraction of the time and memory footprint (see Sec.˜4.3).
  • Efficacy. As shown in Sec.˜5.1, Core Space merging improves performance over full-space merging when non-linear methods are used. In Sec.˜5.2, we show that this improvement stems from better alignment and more compact representation of task-specific directions.
@@ 4.2 No Information Loss in Core Space Representation
Replacing $U_{B}^{(t)}$ and $V_{A}^{(t)}$ with $U_{B}^{\mathrm{ref}}R_{B}^{(t)}$ and $V_{A}^{\mathrm{ref}}Q_{A}^{(t)}$ to obtain Eqs.˜6 and 7, which define the final form of the core matrix, requires that the solutions to the least-squares problems in Eq.˜4 incur zero alignment error. That is,
$\left\|U_{B}^{\mathrm{ref}}R_{B}^{(t)}-U_{B}^{(t)}\right\|_{F}^{2}=0,\quad\quad\left\|{V_{A}^{\mathrm{ref}}}Q_{A}^{(t)}-{V_{A}^{(t)}}\right\|_{F}^{2}=0.$ (12)
In this section, we show that the reference bases $U_{B}^{\mathrm{ref}}$ and $V_{A}^{\mathrm{ref}}$, obtained via the SVD of the stacked matrices $B^{(t)}$ and $A^{(t)}$ (see Eq.˜3), minimize the total alignment error across all $T$ tasks, achieving an error of exactly zero. To illustrate this, we first analyze the alignment error for a single task $t$, focusing on $U_{B}^{\mathrm{ref}}$. Analogous results hold symmetrically for $V_{A}^{\mathrm{ref}}$. For clarity, we assume in the following derivations that $T\cdot r\leq m$ and $T\cdot r\leq n$, so that the total LoRA rank does not exceed the maximum possible rank of the target weight matrix. In Sec.˜A.4, we provide a more general analysis that removes this assumption and demonstrate that the zero alignment error result continues to hold.
@@
$\begin{gathered}\max_{U\in\mathcal{S}}\left\|{U_{B}^{(t)}}^{\top}U\right\|_{F}^{2}=\max_{U\in\mathcal{S}}\mathrm{tr}\left(U^{\top}{U_{B}^{(t)}}{U_{B}^{(t)}}^{\top}U\right),\\ \text{where}\quad\mathcal{S}=\left\{U\in\mathbb{R}^{m\times Tr}\;\middle|\;U^{\top}U=I_{T\cdot r}\right\}\end{gathered}$ (14)
and $\mathrm{tr}(\cdot)$ denotes the trace operator. The optimization domain is restricted to the Stiefel manifold $\mathcal{S}$ (i.e., the set of matrices with orthonormal columns). The following lemma characterizes the solution to the following optimization problem:
@@
A solution $U^{*}$ to the quadratic program in Eq.˜14 is given by a basis whose columns include the $r$ eigenvectors corresponding to nonzero eigenvalues of $B^{(t)}{B^{(t)}}^{\top}\in\mathbb{R}^{m\times m}$ or, equivalently, by the $r$ left singular vectors of the matrix $B^{(t)}$. Moreover, at the optimum, the objective attains its maximum value $r$, resulting in zero alignment error in Eq.˜13.
We refer the reader to Sec.˜A.3 for a detailed proof. Briefly, the result follows by applying the method of Lagrange multipliers to augment the optimization objective with the Stiefel manifold constraint and then enforcing stationarity by setting the gradient of the Lagrangian to zero.
@@
3:Compute $V_{A}^{\mathrm{ref}}$: $\text{stack}(A^{(t)})=U_{A}\Sigma_{A}\underline{V_{A}^{\mathrm{ref}\top}}$ $\triangleright$ reference bases
4:Compute $U_{B}^{\mathrm{ref}}$: $\text{stack}(B^{(t)})=\underline{U_{B}^{\mathrm{ref}}}\Sigma_{B}V_{B}^{\top}$
5:for $t=1$ to $T$ do
6:  Compute: $M^{(t)}=({U_{B}^{\mathrm{ref}}}^{\top}B^{(t)})(A^{(t)}V_{A}^{\mathrm{ref}})$ $\triangleright$ Eq.˜8
7:Merge aligned core matrices: $M_{\mathrm{merged}}=\mathcal{M}(\{M^{(t)}\}_{t=1}^{T})$
8:return $\Delta W=U_{B}^{\mathrm{ref}}M_{\mathrm{merged}}V_{A}^{\mathrm{ref}\top}$ $\triangleright$ reconstructed merged model
Extension to multiple tasks. Achieving zero reconstruction error for a single model $t$ does not guarantee optimality for any other model $t^{\prime}\neq t$. Therefore, we aim to identify a reference basis $U^{*}$ that jointly optimizes Eq.˜14 across all $T$ models. We formulate this global problem as:
where $\mathbf{U}_{B}=\big[U_{B}^{(1)},\;\;U_{B}^{(2)},\;\;\dots,\;\;U_{B}^{(T)}\big]$ denotes the horizontal concatenation of all $U_{B}^{(t)}$ matrices. The equality in Eq.˜15 follows directly from the linearity of the trace operator and the distributivity of matrix multiplication concerning matrix addition: $M^{\top}A_{1}M+M^{\top}A_{2}M=M^{\top}(A_{1}+A_{2})M$.